# What Sensors See, What People Feel: Exploring Subjective Collaboration Perception in Mixed Reality

### v1

@@ 5.1 Implications
Our study indicates key takeaways for designing collaborative MR systems and for future research that aims to bridge sensed behaviors with user experiences. The link between shared gaze and perceived collaboration implies that visual coordination acts as a behavioral signal and a perceptual anchor for participants. Therefore, MR systems should feature indicators such as shared gaze highlights or real-time attention cues to enhance awareness of group focus. Such simple interventions can foster shared attention and keep users in sync without disrupting task momentum.
Across multiple hypotheses, group-level behaviors more accurately predicted collaboration and task performance than individual metrics. This suggests that meaningful social patterns emerge through aggregation, not simply from individual activity. Systems that infer group dynamics or adapt to team behavior should avoid user-centric interpretations in isolation. Instead, they should incorporate group-level features that reflect distribution, synchronicity, and co-occurrence.
Our results show that while sensed behavior reliably maps onto collaborative perception and task efficiency, it falls short of capturing deeper subjective experiences, such as presence or cognitive availability in its current form. This reinforces that not all internal states are externally observable, even in high-fidelity sensing environments. To meaningfully comprehend experience from sensor data, some intermediary techniques are needed to connect high-level behavioral indicators with low-level experiential states. Designers should be cautious about over-interpreting behavioral signals as proxies for mental state and instead consider combining sensing with self-reports or adaptive prompting. Hybrid models that integrate system-logged data with lightweight subjective inputs could enable more robust personalization and responsiveness in collaborative MR.
@@ 6 Limitations and Future Work
Group composition was not experimentally manipulated; participants had the autonomy to self-select or be randomly assigned to groups. This decision sought to observe genuine interaction dynamics instead of creating artificial teams. Nonetheless, pre-existing familiarity among participants might impact conversation, shared focus, and comfort. Although not measured in this study, future research could quantify prior familiarity to understand its effects better.
In addition, we focus on interaction signals such as gaze, proximity, and voice activity, which could be tracked using standard MR hardware. This choice aligned with our aim to create practical, lightweight sensing systems. Yet, using only headset sensors restricted us from accessing detailed affective, physiological, or facial data. Future research should investigate multimodal enhancements for richer sensing, such as emotion recognition or biometric monitoring.
Finally, while our framework bridges sensor data and subjective reflection, it cannot fully capture internal states such as frustration, motivation, or attention lapses. These nuanced experiences often remain invisible to behavioral sensing alone. Our study highlights the gap between observable behavior and internal experience, reinforcing the need for interpretive models that integrate multiple modalities and participant feedback.
@@ Acknowledgements.
This work is supported by the U.S. National Science Foundation (NSF) under grant number 2339266 and 2237485.References
@@
- [4] H. Bai, P. Sasikumar, J. Yang, and M. Billinghurst. A user study on mixed reality remote collaboration with eye gaze and hand gesture sharing. In Proceedings of the ACM Conference on Human Factors in Computing Systems, 2020.
@@
- [9] M. D. Bonis, H. Nguyen, and P. Bourdot. A literature review of user studies in extended reality applications for archaeology. In 2022 IEEE International Symposium on Mixed and Augmented Reality, pp. 92–101, 2022.
@@
- [26] H. Le Tarnec, O. Augereau, E. Bevacqua, and P. De Loor. Improving Collaborative Learning in Virtual Reality With Facial Expressions. In Adjunct Proceedings of the 2022 ACM International Joint Conference on Pervasive and Ubiquitous Computing and the 2022 ACM International Symposium on Wearable Computers, UbiComp/ISWC ’22 Adjunct. New York, NY, USA, 2023.
@@
- [38] I. R. Pinto, J. M. Marques, J. M. Levine, and D. Abrams. Membership status and subjective group dynamics: Who triggers the black sheep effect? Journal of personality and social psychology, 99(1):107, 2010.
@@
- [42] T. Postmes, R. Spears, K. Sakhel, and D. De Groot. Social influence in computer-mediated communication: The effects of anonymity on group behavior. Personality and Social Psychology Bulletin, 27(10), 2001.
@@
- [45] J. A. Russell. A circumplex model of affect. Journal of Personality and Social Psychology, 39(6), 1980.
@@
- [64] X. Xu, A. Puggioni, D. Kilroy, and A. G. Campbell. User experience of collaborative co-located mixed reality: a user study in teaching veterinary radiation safety rules. In IEEE International Symposium on Mixed and Augmented Reality (ISMAR), 2023.

### v2

@@ 5.1 Implications
Our study indicates three key takeaways for designing collaborative MR systems and for future research that aims to bridge sensed behaviors with user experiences. The link between shared gaze and perceived collaboration implies that visual coordination acts as a behavioral signal and a perceptual anchor for participants (1). Therefore, MR systems should feature indicators such as shared gaze highlights or real-time attention cues to keep users in sync without disrupting task momentum.
Across multiple hypotheses, group-level behaviors more accurately predicted collaboration and task performance than individual metrics. This suggests that meaningful social patterns emerge through aggregation (2), not simply from individual activity. Systems that adapt to team behavior should avoid user-centric interpretations in isolation. Instead, they should incorporate group-level features that reflect distribution, synchronicity, and co-occurrence.
Our results show that while sensed behavior reliably maps onto collaborative perception and task efficiency, it falls short of capturing deeper subjective experiences, such as presence or cognitive availability, in its current form. This reinforces that not all internal states are externally observable (3), even in high-fidelity sensing environments. This absence of mapping is not a flaw it is a vital empirical finding that delineates what current MR sensing can (and importantly cannot) reveal about subjective experience. While prior works have treated gaze or physiological synchrony as proxies for presence or co-presence [28], our results underscore their limitations without deeper context. Our findings align with recent studies showing that even rich physiological or behavioral data often fail to reliably predict internal states unless complemented with richer modalities, such as neural, situational, or self-report inputs [40, 25, 29]. To meaningfully comprehend experience from sensor data, some intermediary techniques are needed to connect high-level behavioral indicators with low-level experiential states. Designers should treat behavioral signals as interpretive anchors, not definitive measures of internal states, reinforcing the cautious approach seen in immersive cognitive sensing research, and instead consider combining sensing with self-reports or adaptive prompting. Hybrid models that integrate system-logged data with lightweight subjective inputs could enable more robust personalization and responsiveness in collaborative MR.
@@ 6 Limitations and Future Work
Group composition was not experimentally manipulated; participants had the autonomy to self-select or be randomly assigned to groups. This decision sought to observe genuine interaction dynamics instead of creating artificial teams. Nonetheless, pre-existing familiarity among participants might impact conversation, shared focus, and comfort. Although not measured in this study, future research could quantify prior familiarity to better understand its effects.
In addition, we focus on interaction signals such as gaze, proximity, and voice activity, which could be tracked using standard MR hardware. This choice aligned with our aim to create practical, lightweight sensing systems. Yet, using only headset sensors restricted us from accessing detailed affective, physiological, or facial data. Future research should investigate multimodal enhancements for richer sensing, such as emotion recognition or biometric monitoring. Finally, while our framework bridges sensor data and subjective reflection, it cannot fully capture internal states such as frustration, motivation, or attention lapses. These nuanced experiences often remain invisible to behavioral sensing alone. Our study highlights the gap between observable behavior and internal experience, reinforcing the need for interpretive models that integrate multiple modalities and participant feedback.
@@ Acknowledgements.
This work is supported by the U.S. National Science Foundation (NSF) under grant numbers 2339266 and 2237485.References
@@
- [4] H. Bai, P. Sasikumar, J. Yang, and M. Billinghurst. A user study on mixed reality remote collaboration with eye gaze and hand gesture sharing. In Proceedings of the 2020 CHI conference on human factors in computing systems, pp. 1–13, 2020.
- [5] J. N. Bailenson, A. C. Beall, and J. Blascovich. Gaze and task performance in shared virtual environments. The Journal of Visualization and Computer Animation, 13(5):313–320, 2002. _eprint: https://onlinelibrary.wiley.com/doi/pdf/10.1002/vis.297. doi: 10 . 1002/vis . 297
@@
- [9] A. Bayro, Y. Ghasemi, and H. Jeong. Subjective and objective analyses of collaboration and co-presence in a virtual reality remote environment. In 2022 IEEE Conference on Virtual Reality and 3D User Interfaces Abstracts and Workshops (VRW). IEEE, 2022.
@@
- [11] D. Bohus and E. Horvitz. Facilitating multiparty dialog with gaze, gesture, and speech. In International Conference on Multimodal Interfaces and the Workshop on Machine Learning for Multimodal Interaction, ICMI-MLMI ’10. Association for Computing Machinery, New York, NY, USA, 2010. doi: 10 . 1145/1891903 . 1891910
@@
- [17] R. A. J. de Belen, H. Nguyen, D. Filonik, D. Del Favero, and T. Bednarz. A systematic review of the current state of collaborative mixed reality technologies: 2013–2018. AIMS Electronics and Electrical Engineering, 3(2):181–223, 2019.
@@
- [25] Y. Hou, Q. Xie, N. Zhang, and J. Lv. Cognitive load classification of mixed reality human computer interaction tasks based on multimodal sensor signals. Scientific Reports, 15(1):13732, 2025.
@@
- [28] A. Jing, K. W. May, M. Naeem, G. Lee, and M. Billinghurst. eyemr-vis: Using bi-directional gaze behavioural cues to improve mixed reality remote collaboration. In Extended Abstracts of the 2021 CHI Conference on Human Factors in Computing Systems, CHI EA ’21. Association for Computing Machinery, New York, NY, USA, 2021. doi: 10 . 1145/3411763 . 3451844
- [29] G. Kim, J. Lee, D. Yeo, E. An, and S. Kim. Physiological indices to predict driver situation awareness in vr. In Adjunct Proceedings of the 2023 ACM International Joint Conference on Pervasive and Ubiquitous Computing & the 2023 ACM International Symposium on Wearable Computing, pp. 40–45, 2023.
@@
- [48] M. A. Rosen, A. S. Dietz, T. Yang, C. E. Priebe, and P. J. Pronovost. An integrative framework for sensor-based measurement of teamwork in healthcare. Journal of the American Medical Informatics Association, 22(1):11–18, 2015.
- [49] J. A. Russell. A circumplex model of affect. Journal of Personality and Social Psychology, 39(6):1161–1178, Dec. 1980. doi: 10 . 1037/h0077714
@@
- [60] J. Thomas, S. W. Lee, A. Giovannelli, L. Lane, and D. Bowman. A communication-focused framework for understanding immersive collaboration experiences. In 2023 IEEE Conference on Virtual Reality and 3D User Interfaces Abstracts and Workshops (VRW), pp. 301–304. IEEE, 2023.
- [61] L. Thomaschewski, N. Feld, B. Weyers, and A. Kluge. I sense that there is someone else: an exploratory study on the influence of the media richness of augmented reality-based assistance systems on team experience and performance. Frontiers in Virtual Reality, 4, 2023.